# Generalization performance of narrow one-hidden layer networks in the teacher-student setting

### v3

@@ 4 Discussion and conclusion
We developed a theory for the typical generalization performance of a narrow one-hidden-layer network in the teacher–student setting, under both finite-temperature Bayesian learning and empirical risk minimization. Our theory highlights the presence of distinct phases and thermodynamic states, including metastable ones that can trap the dynamics of full-batch GD, both with and without explicit noise. We further showed that the activation function is responsible for the appearance of different types of transitions in a Bayesian setting, in a manner that depends on the the temperature parameter. Interestingly, we identified a memorization phase in which the student network breaks permutation symmetry among its hidden units without aligning to the teacher’s weights. This study differs from the recent work of [oostwal_hidden_2021, citton_phase_2025], which focuses on the large-temperature, infinite-data regime, and from [nishiyama_solution_2025], which investigates the storage capacity problem. It also goes beyond classical committee-machine analysis [schwarze_statistical_1992, schwarze_learning_1993, engel_storage_1992, schwarze_discontinuous_1993, schwarze_learning_1993-1, urbanczik1997full, ahr1999statistical] by offering a theoretical framework valid for arbitrary inner and outer activation functions with regularized weights.
@@
Learning in two-layer networks in the $K\ll N$ regime has been recently analyzed using dynamical mean field theory for gradient flow [montanari_dfmt_2025]. The authors consider both first- and second-layer learnable weights and identify different dynamical regimes using a separation of time-scales argument. It would be interesting to characterize the steady states of such learning dynamics, in relation to the equilibrium distribution induced by Langevin-based sampling of the weight posterior. Doing so would require extending our framework to the case of plastic second-layer weights, which we treated as known (quenched) for simplicity. This extension may also help in fully clarifying the difference between the $K\ll N$ case and the so-called extensive-width regime, i.e. with $K=O\left(N\right)$ and $P=O\left(N\right)$, see [pacelli2023statistical] and references therein.
@@
Generalization performance of narrow one-hidden layer networks in the teacher-student setting
@@ Appendix B The limit of large number of hidden units
The replica analysis led to a major simplification: it allowed us to transform the original optimization problem in eq.(20) spanning over an $N\times K$-dimensional space, into the $K\times K$-dimensional extremization problem in eq. (50). Solving this problem analytically for an arbitrary $K$ is generally intractable. In this work, we consider the limit where $K\rightarrow\infty$ but slower than $N$, meaning that $K/N\rightarrow 0$ as in classical and recent works on committee machines [barkai_broken_1992, engel_storage_1992, baldassi_properties_2019, schwarze_discontinuous_1993, schwarze_learning_1993-1, aubin_committee_2018, nishiyama_solution_2025]. In this limit, as suggested in [schwarze_learning_1993], it is reasonable to look for solutions of eq. (50) of the form:
@@ B.2 Energetic potential
This Kernel function is called Neural Network Gaussian Process (NNGP) and describes the covariance of the function implemented by a neural network at initialization (i.e., with random weights) in the infinite-width limit, evaluated at two different inputs [Neal1996, Williams1996].
@@ B.2.2 Classification case
Note that this expression reduces to the ones found by Schwarze [schwarze_learning_1993] in the case of sign activation function case, zero label noise $\Delta^{\star}$ and theta loss $\ell(x)=\Theta(-x)$ in the infinite $\beta$ limit. Similarly, this expression reduces to the one reported in a recent paper [nishiyama_solution_2025] studying the storage problem, where the labels are extracted completely random. This corresponds to the limit $\Delta^{\star}\to\infty$ of our expressions, where in addition has $m_{a}=m_{d}=0$.
@@ E.3 Bayes optimal case
In the Bayes optimal case the saddle point equations are particularly simply to write in both the small and large data regimes described above. Indeed, due to matching of the teacher and student prior and likelihood, one has the following Nishimori conditions [Nishimori_1980]

### v4

@@ 5 Discussion and conclusion
We developed a theory for the typical generalization performance of a narrow one-hidden-layer network in the teacher–student setting, under both finite-temperature Bayesian learning and empirical risk minimization. Our theory highlights the presence of distinct phases and thermodynamic states, including metastable ones that can trap the dynamics of full-batch GD, both with and without explicit noise. We further showed that the activation function is responsible for the appearance of different types of transitions in a Bayesian setting, in a manner that depends on the the temperature parameter. Interestingly, we identified a memorization phase in which the student network breaks permutation symmetry among its hidden units without aligning to the teacher’s weights. This study differs from the recent work of [35, 13], which focuses on the large-temperature, infinite-data regime, and from [34], which investigates the storage capacity problem. It also goes beyond classical committee-machine analysis [44, 43, 15, 45, 42, 51, 3] by offering a theoretical framework valid for arbitrary inner and outer activation functions with regularized weights.
@@
Learning in two-layer networks in the $K\ll N$ regime has been recently analyzed using dynamical mean field theory for gradient flow [30]. The authors consider both first- and second-layer learnable weights and identify different dynamical regimes using a separation of time-scales argument. It would be interesting to characterize the steady states of such learning dynamics, in relation to the equilibrium distribution induced by Langevin-based sampling of the weight posterior. Doing so would require extending our framework to the case of plastic second-layer weights, which we treated as known (quenched) for simplicity. This extension also help in fully clarifying the difference between the $K\ll N$ case and the so-called extensive-width regime, i.e. with $K=O\left(N\right)$ and $P=O\left(N\right)$, see [24, 36] and references therein.
@@ Appendix B The limit of large number of hidden units
The replica analysis led to a major simplification: it allowed us to transform the original optimization problem in eq.(20) spanning over an $N\times K$-dimensional space, into the $K\times K$-dimensional extremization problem in eq. (50). Solving this problem analytically for an arbitrary $K$ is generally intractable. In this work, we consider the limit where $K\rightarrow\infty$ but slower than $N$, meaning that $K/N\rightarrow 0$ as in classical and recent works on committee machines [10, 15, 7, 45, 42, 6, 34]. In this limit, as suggested in [43], it is reasonable to look for solutions of eq. (50) of the form:
@@ B.2 Energetic potential
This Kernel function is called Neural Network Gaussian Process (NNGP) and describes the covariance of the function implemented by a neural network at initialization (i.e., with random weights) in the infinite-width limit, evaluated at two different inputs [32, 52].
@@ B.2.2 Classification case
Note that this expression reduces to the ones found by Schwarze [43] in the case of sign activation function case, zero label noise $\Delta^{\star}$ and theta loss $\ell(x)=\Theta(-x)$ in the infinite $\beta$ limit. Similarly, this expression reduces to the one reported in a recent paper [34] studying the storage problem, where the labels are extracted completely random. This corresponds to the limit $\Delta^{\star}\to\infty$ of our expressions, where in addition has $m_{a}=m_{d}=0$.
@@ E.3 Bayes optimal case
In the Bayes optimal case the saddle point equations are particularly simply to write in both the small and large data regimes described above. Indeed, due to matching of the teacher and student prior and likelihood, one has the following Nishimori conditions [33]